OpenStudy (anonymous):

Solving with the interval [0,2pi) 1+sin(theta)=1/2

OpenStudy (anonymous):

x= (7/6)pi, and x= (11/6)pi ?

OpenStudy (anonymous):

\[\sin \theta=(1/2)-1\] \[\sin \theta=-1/2\] If you have the interval [0,2pi], then you are looking for an angle (or angles) between 0 and 2pi whose sin equals -1/2

OpenStudy (anonymous):

yes, that would be the answer for your question.

OpenStudy (anonymous):

Ah thanks